OpenStudy (anonymous):

What is a possible first step to eliminate the variable x in the following system? 2x – 7y = 3 4x + 5y = –2

OpenStudy (anonymous):

multiply the first equation by -2. to make it -4x+14y=-6. cancel out variable x after.

OpenStudy (anonymous):

divide the second by -2

OpenStudy (anonymous):

multiply the 1st equation on -2 then add both equations

OpenStudy (anonymous):

2x – 7y = 3 4x + 5y = –2 (2x – 7y = 3)*4 (4x + 5y = –2)*2 8x-28y=12------->1 8x+20y=-8------->2 subtract 2 from 1 (8x-28y=12)-(8x+20y=-8) 8x-28y=12 -8x-20y=8 0-48y=20 24y=10 y=5/12